OpenStudy (anonymous):

Describe the concavity of the graph of f on the interval [0, π]. f(x)=3x^2-3sin(2x)

OpenStudy (anonymous):

You need to find \[f\prime \prime(x)\] and look at this function over the range [0,pi] when the function is positive/negative it means positive/negative concavity

OpenStudy (anonymous):

I know the derivative of it is 6x-6cos2x but when I equal it to zero \[x-\cos2x=0\]

OpenStudy (anonymous):

Oh so I dont need to pay attention for the first derivative

OpenStudy (anonymous):

No not for concavity. First derivative test is to find the extreme points of f(x)

OpenStudy (anonymous):

Oh ok then the second derivative is \[6+12\sin(x)\] So then I equal that to zero correct?

OpenStudy (anonymous):

I ment the second derivative is \[6+12\sin(2x)\]

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

Yes set to zero and look for positive and negative values

OpenStudy (anonymous):

So far I got \[\sin(2x)=\frac{ -1 }{ 2 }\]

OpenStudy (anonymous):

No sorry not set to zero... you plug in the domain [0, pi] and see when it is positive or negative

OpenStudy (anonymous):

ok then its positive from 0 to pi/2 then negetive at 2pi/3 to 5pi/6 then positive at pi

OpenStudy (anonymous):

But both pi and 0 are positive

OpenStudy (anonymous):

Im confused because my awnser choices are a)concave up on (0, 7π⁄12 ); concave down on ( 7π⁄12, π). b)concave up on ( 7π⁄12, 11π⁄12 ); concave down on (0, 7π⁄12 ) and on ( 11π⁄12, π). c)concave down on (0, π). d)concave up on (0, 7π⁄12 ) and on ( 11π⁄12, π); concave down on ( 7π⁄12, 11π⁄12 ). e)concave up on (0, π⁄12 ); concave down on ( π⁄12, π).

OpenStudy (anonymous):

I graphed the function and found positive (concave up) from (0, 7pi/12) negative (concave down) from (7pi/12, 11pi/12) positive (concave up) from (11pi/12, pi)

OpenStudy (anonymous):

I think you were missing some number when you performed the last step

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6%2B12sin%282x%29 You can see a graph here of the function

OpenStudy (anonymous):

I guess if I needed to do it without a calculator I needed to solve the second derivative

OpenStudy (anonymous):

and set it as equal to zero correct?

OpenStudy (anonymous):

No you dont need to set the second derivative to zero because that does not tell you any information. Only if the second derivative is positive or negative does it tell you the concavity. You set the first derivative equal to zero to find the extreme points of the original function.

OpenStudy (anonymous):

Another type of question with this is to find the extreme points of f(x) with the first derivative test and then use the second derivative at those points to see if they are a max( negative concavity) or a minimum (positive concavity)